### v1

# Manual-PA: Learning 3D Part Assembly from Instruction Diagrams

Abstract
Assembling furniture amounts to solving the discrete-continuous optimization task of selecting the furniture parts to assemble and estimating their connecting poses in a physically realistic manner. The problem is hampered by its combinatorially large yet sparse solution space thus making learning to assemble a challenging task for current machine learning models. In this paper, we attempt to solve this task by leveraging the assembly instructions provided in diagrammatic manuals that typically accompany the furniture parts. Our key insight is to use the cues in these diagrams to split the problem into discrete and continuous phases. Specifically, we present Manual-PA, a transformer-based instruction Manual-guided 3D Part Assembly framework that learns to semantically align 3D parts with their illustrations in the manuals using a contrastive learning backbone towards predicting the assembly order and infers the 6D pose of each part via relating it to the final furniture depicted in the manual. To validate the efficacy of our method, we conduct experiments on the benchmark PartNet dataset. Our results show that using the diagrams and the order of the parts lead to significant improvements in assembly performance against the state of the art. Further, Manual-PA demonstrates strong generalization to real-world IKEA furniture assembly on the IKEA-Manual dataset.
1 Introduction
Figure 1: An illustration of the manual-guided 3D part assembly task. Given (a) a diagrammatic manual book demonstrating the step-by-step assembly process and (b) a set of texture-less furniture parts, the goal is to (c) infer the order of parts for the assembly from the manual sequence and predict the 6DoF pose for each part such that the spatially transformed parts assembles the furniture described in the manual.
With the rise of the DIY trend, nowadays it is the consumers who are increasingly assembling products like IKEA furniture [20, 2, 53, 54, 3], LEGO structures [46, 44], and exercise equipments [57, 42]. This paradigm shift can reduce the product cost, but often poses challenges for untrained consumers, as an incorrect assembly in any step may lead to irreparable outcomes. Thus, automating the assembly or providing active help in the process is often desired.
Automated assembly faces the dual challenge of large combinatorial solution space, parameterized by the number of discrete parts in the assembly, and the need for a precise 6D pose estimation that correctly connects parts between each other. The limited number of feasible, stable assembly sequences makes this discrete-continuous optimization problem extremely challenging, even for advanced neural networks, especially as the number of parts increases [51]. Fortunately, most assembly tasks include language-agnostic instruction manuals, such as diagrams, delineating the assembly. While such manuals help shrink the solution search space, the problem of correctly matching the 2D manual diagrams with their corresponding 3D object parts and inferring their assembly pose and order remains a significant challenge for automated systems.
This paper introduces a method to enable machines to learn how to assemble shapes by following a visual instruction manual, a task we call manual-guided 3D part assembly. As illustrated in Fig. 1, given a set of parts and a manual, our approach interprets the manual’s step-by-step diagrams to predict the precise 6D pose of each part, assembling them into the target shape. There are two lines of prior approaches that have similarities to our proposed task, namely methods that focus on the geometry of individual parts and inter-part relationships [22, 52, 55, 8, 56, 9], and methods such as MEPNet [46] that learns the assembly in specific conditions, such as LEGO shapes. The former rely on priors like peg-hole joints [24] or sequence generation [51] to reduce search complexity but are generative and may yield unstable assemblies due to occlusions [22]. The latter often simplifies the task by assuming that parts are provided step-by-step, e.g., LEGO manuals. However, furniture manuals, such as IKEA, may not have such information, and it is necessary to determine which parts are used at each step before assembly. Additionally, while LEGO tasks feature standardized “stud” joints, general furniture assembly is more complex and lacks such constraints.
Based on the above observations, we identify two key challenges in our task. First, how to detect which part or parts are added at each step in the assembly diagrams. Given that these diagrams depict incremental assembly stages, the core challenge is to recognize newly added parts at each step. This requires aligning 2D assembly diagrams with the 3D parts to establish a correct assembly sequence. Second, how to effectively align the assembly process with the learned order. A straightforward approach is to directly use the learned sequence as the order for predicting poses, e.g., by adopting an auto-regressive prediction method similar to MEPNet [46]. However, this approach carries the risk of accumulating errors—if a mistake occurs in an earlier step, subsequent steps are likely to fail as well. On the other hand, completely disregarding the sequence forces the assembly model to implicitly find correspondences, which makes both learning and prediction significantly more difficult. Therefore, the inferred sequence should serve as a soft guidance for the assembly process, helping each part focus more on the relevant step diagram.
To address these challenges, we present Manual-guided Part Assembly (Manual-PA), a novel transformer-based neural network that predicts correspondences between step diagrams and parts by computing their semantic similarity, which is learned using contrastive learning. Using this similarity matrix, we establish an assembly order via solving an optimal assignment problem between the aligned features of the 3D object parts and the ordered manual diagram steps, followed by permuting each part’s positional encoding according to the learned order. Finally, a transformer decoder fuses the two feature modalities to predict the final 6DoF pose for each part. Notably, the cross-attention between step diagrams and parts is guided by the positional encoding, ensuring that each step diagram receives a higher attention score for its corresponding part.
To empirically validate the effectiveness of Manual-PA, we conduct experiments on the standard PartNet benchmark dataset [26], on the categories of assembling the Chair and Table classes. Our experiments and ablation studies clearly show that Manual-PA leads to state-of-the-art results on multiple metrics, including success rate, outperforming prior methods by a significant margin. Results further showcase the strong generalizability of our approach to real-world IKEA furniture assembly through experiments using the IKEA-Manual dataset [47].
We summarize our key contributions below:
  1. 1.
    We introduce a new problem setup of assembling 3D parts using diagrammatic manuals for guidance.
  2. 2.
    We present a novel framework, Manual-PA, for solving this task that learns the assembly order of the 3D parts, which is then used as soft guidance for the assembly.
  3. 3.
    We present experiments on the benchmark PartNet and IKEA-Manual datasets demonstrating state-of-the-art performances.
2 Related Works
2.0.1 3D Part Assembly
involves reconstructing complete shapes from sets of 3D parts, a task originally assisted by leveraging semantic part information [6, 14, 49]. With the introduction of the PartNet dataset [26], recent works [52, 27, 8, 55, 9, 56] leverage geometric features. For instance, 3DHPA [9] uses part symmetry for a hierarchical assembly, SPAFormer [51] circumvents combinatorial explosion by generating a part order, and Joint-PA [24] introduces a peg-hole constraint approach to facilitate assembly. However, these methods are generative towards producing plausible final shapes without aiming for preciseness. Approaches like GPAT [23] and Image-PA [22] add guidance from point clouds or rendered images, respectively. Our work extends this direction by introducing guidance through instruction manuals, similar to how humans would do.
2.0.2 Permutation Learning
is often used in self-supervised methods for learning image representations [30, 39, 15, 4, 34]. However, in this work, we seek to align the 3D parts to instruction steps, for which we adopt a contrastive learning  [16, 13, 21, 38] using a novel cross-modal alignment setup.
2.0.3 6DoF Object Pose Estimation
methods can be classified into instance-level [37, 35, 50], category-level [41, 45], and unseen [19, 48]. Our approach relates to the unseen object category as furniture typically comprises novel parts. When the 3D models of these parts are available, many methods [7, 33, 29, 25, 12] rely on feature matching followed by the PnP algorithm to infer pose. However, IKEA-style manuals are represented as line drawings making accurate feature extraction for matching to 3D parts challenging. Alternatively, some approaches [19, 48] use end-to-end models to directly predict the 6D pose. A common approach is to use CNOS [28] for object segmentation. However, in assembly diagrams, where multiple parts are connected and depicted together, SAM-based methods [17] often struggle with accurate segmentation. Our approach addresses these challenges using an end-to-end framework that uses a cross-attention scheme to enable each part to attend to its relevant region in the step diagram.
Figure 2: Overview of our proposed method Manual-PA. (a) Feature extraction (Sec. 3.2): we extract semantic and geometrical features from both the step diagrams of the assembly manual and the corresponding part point clouds using the image encoder and point encoder, respectively. (b) Manual-guided part permutation learning (Sec. 3.3): we compute a similarity matrix $\mathbf{S}$ between the two modalities, and subsequently apply the Hungarian algorithm to obtain the permutation matrix $\mathbf{P}$ for the parts. (c) Manual-guided part pose estimation (Sec. 3.4): we add positional encodings (PE) $\Phi$ to both the step diagrams and parts, where the PE order for parts is determined by the order predictions from (b). This is followed by a transformer decoder to enable multimodal feature fusion and interaction, along with a pose prediction head to determine the rotation $R_{i}$ and translation $t_{i}$ of each part. The predicted poses are then applied to the corresponding parts to obtain the final assembled shape $\mathbf{S}$.
3 Manual-Guided Part Assembly
In this section, we first provide a concrete formal definition of the manual-based 3D part assembly task in Sec. 3.1. We then present a detailed explanation in Secs. 3.2, 3.3 and 3.4 of the proposed learning pipeline as depicted in Fig. 2, followed by a description of the training process and the loss functions used for each stage in Sec. 3.5.
3.1 Problem Formulation
Our task involves two different input modalities: a set of 3D parts and a corresponding instruction manual. The 3D parts are represented as a multiset [^1] We allow multiple identical parts, e.g., chair legs, to appear in the set. of unordered point clouds, denoted by $\{\mathcal{P}_{i}\}_{i=1}^{N}$. Note that the number of parts $N$ may differ from furniture to furniture. The manual consists of an ordered sequence of $N$ step diagram images, denoted by $(\mathcal{I}_{1},\mathcal{I}_{2},\ldots,\mathcal{I}_{N})$; each diagram depicting an incremental step in the assembly process of a 3D furniture shape $\mathcal{S}$. These step diagrams are 2D line drawings, which are 2D projections of textureless 3D CAD models of these parts captured from a camera $\mathcal{C}$, and we assume for each step, only one additional part is attached to the furniture assembled thus far. Our goal is to align the 3D parts with their delineations in the respective manual steps towards predicting the assembly order of the parts and the part poses $\{(R_{i},t_{i})\in SE(3)\}_{i=1}^{N}$ for the final assembled furniture, where $R_{i},t_{i}$ represent the $3\times 3$ rotation matrix and the three-dimensional translation vector respectively for the $i$-th part in the SE(3) transformation space. That is, these parts undergo rigid transformations such that their combination constructs the complete 3D furniture shape $\mathcal{S}$ as seen from the viewpoint of camera $\mathcal{C}$.
3.2 Feature Extraction
We first use off-the-shelf encoders to extract semantic latent features from the inputs. For the 3D part point clouds $\{\mathcal{P}_{i}\}_{i=1}^{N}$, we utilize a lightweight version of PointNet [36], similar to the one used in Image-PA [22], followed by a linear layer to map the features to a common dimension $D$, resulting in geometric part features ${\mathbf{f}^{\text{P}}}\in\mathbb{R}^{N\times D}$. For sequences of step diagrams $(\mathcal{I}_{j})_{j=1}^{N}$, given that the assembly instructions are provided step-by-step in an incremental manner, our primary focus is on the differences between consecutive steps. For any two consecutive diagram images $\mathcal{I}_{j}$ and $\mathcal{I}_{j+1}$, for $j=1,2,\ldots,N-1$, we compute the difference image given by $|\mathcal{I}_{j}-\mathcal{I}_{j+1}|$ that shows which new part is added to the thus-far assembled furniture. Next, this difference image is patchified into $K$ image patches, that are then fed to DINOv2 image encoder [5, 32] to extract features, followed by a linear layer to map these features to the same dimensionality $D$, resulting in semantic image features ${\mathbf{f}^{\text{I}}}\in\mathbb{R}^{N\times K\times D}$.
3.3 Manual-Guided Part Permutation Learning
In our task, we seek to learn the order of a set of parts conditioned on an ordered sequence of step diagrams. This problem is similar to learning the feature correspondences between two modalities in order to align them, with an ideal alignment corresponding to a permutation matrix. Specifically, our objective is to derive a similarity matrix between the two modalities such that each modality feature is unambiguously assigned to its corresponding feature in the other modality and we desire this assignment matches with the ground truth order in the step sequence in the manual. Formally, we construct the similarity matrix $\mathbf{S}\in\mathbb{R}^{N\times N}$ as:
$\mathbf{S}_{ij}=\operatorname{sim}({\mathbf{f}_{i}^{\text{P}}},\textbf{g}^{\text{I}}_{j}),$ (1)
where $\textbf{g}^{\text{I}}\in\mathbb{R}^{N\times D}$ is derived from the $N\times K\times D$-dimensional step-diagram features ${\mathbf{f}^{\text{I}}}$ by performing a max-pooling operation on the patch dimension $K$. We measure the similarity in the feature space using the dot product as $\operatorname{sim}(\mathbf{a},\mathbf{b})=\mathbf{a}\cdot\mathbf{b}$.
Next, we obtain the permutation matrix $\mathbf{P}\in\{0,1\}^{N\times N}$ by solving a bipartite matching optimization problem using the Hungarian matching algorithm [18] with cost $\mathbf{C}=-\mathbf{S}$:
$\begin{array}[]{rl}\text{minimize}&\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbf{C}_{ij}\mathbf{P}_{ij}\\ \text{subject to}&\sum_{j=1}^{N}P_{ij}=1,\quad\forall i\in\{1,\dots,N\}\\ &\sum_{i=1}^{N}P_{ij}=1,\quad\forall j\in\{1,\dots,N\}\\ &P_{ij}\in\{0,1\},\quad\forall i,j\in\{1,\dots,N\},\end{array}$ (6)
where we minimize the cost, subject to constraints ensuring $\mathbf{P}$ is a permutation matrix, i.e., each row and column of $\mathbf{P}$ contains exactly one entry equal to $1$, with all other entries being 00. The final part order can be obtained by $\sigma=\operatorname{argmax}(\mathbf{P})$ where $\operatorname{argmax}$ is applied columnwise.
3.4 Manual-Guided Part Pose Estimation
3.4.1 Positional Encoding.
Since the attention mechanism in the transformer is permutation-invariant, we follow Vaswani et al. [43] in using positional encoding to convey the order of both step diagrams and parts to the model. Formally, the positional encoding is defined as a function that maps a scalar to a sinusoidal encoding, $\phi:\mathbb{N}\rightarrow\mathbb{R}^{D}$:
$\phi(x)_{2i}=\sin\left(\frac{x}{\tau_{p}^{2i/d}}\right)\!,\;\phi(x)_{2i-1}=\cos\left(\frac{x}{\tau_{p}^{2i/d}}\right)\!,$ (7)
where $i=1,\dots,d/2$ indexes the dimension and $\tau_{p}$ is a temperature hyperparameter. By applying $\phi$ to encode the sequential order of step diagrams as they appear in the manual (i.e., using $x$ as a simple increasing sequence $1,2,\dots,N$), we obtain a positional encoding $\Phi\in\mathbb{R}^{N\times D}$, where ${\mathbf{p}^{\text{I}}}:=\Phi$. Given the permutation matrix $\mathbf{P}$ that corresponds to the order of the parts, we can compute the positional encoding for the parts as ${\mathbf{p}^{\text{P}}}=\mathbf{P}^{T}\Phi$. During training, we use the ground truth part order, whereas at inference time, we replace it with the predicted permutation matrix $\hat{\mathbf{P}}$, yielding $\hat{\mathbf{p}}^{\text{P}}=\hat{\mathbf{P}}^{T}\Phi$. These positional encodings are then added to the respective features before being fed into the transformer decoder and provide soft guidance on the sequence order.
3.4.2 Transformer Decoder.
The modality interaction and fusion between step diagrams and parts occur through a stack of $L$ transformer [43] decoder layers. In each layer, the self-attention module receives the output from the previous layer (initially, the part features ${\mathbf{f}^{\text{P}}}+{\mathbf{p}^{\text{P}}}$ in the first layer) to perform part-to-part interactions. This is followed by a cross-attention module that injects and fuses the step diagram features ${\mathbf{f}^{\text{I}}}+{\mathbf{p}^{\text{I}}}$ with the part features processed by the self-attention module, enabling step-to-part interactions. Notably, for the step diagram features, we concatenate the features from all diagrams from the same manual along the patch dimension, resulting in a tensor of shape $\mathbb{R}^{NK\times D}$.
3.4.3 Pose Prediction Head.
Finally, the output $O_{\text{Dec}}$ from the transformer decoder is fed into a pose prediction head, which predicts the translation and rotation for each part, denoted as $(\hat{R}_{i},\hat{t}_{i})$. The rotation is represented as a unit quaternion $\hat{q}_{i}\in\mathbb{R}^{4}$. We initialize its bias to an identity quaternion $(1,0,0,0)$ and normalize the prediction to ensure a unit norm.
3.5 Training and Losses
First, we train the manual-guided part order learning model until it successfully aligns the two modalities and provides an optimal permutation for the parts. Subsequently, we train the transformer decoder-based part pose estimation model using the predicted order from the first model to infer the part poses. Our empirical results indicate that having an accurate part order is crucial for enhancing the inference performance of the pose estimation network.
3.5.1 Loss for Permutation Learning.
Borrowing the notation from Secs. 3.2 and 3.3, we define the contrastive-based order loss as:
$\mathcal{L}_{\text{order}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\operatorname{sim}({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{i})/\tau)}{\sum_{j=1}^{B}\exp(\operatorname{sim}({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{j})/\tau)},$ (8)
where $B$ is the mini-batch size, $\tau$ is the temperature, and $\sigma(i)$ denotes the index for the $\sigma(i)$-th part that corresponds to the $j$-th step according to the ground truth part order $\sigma$. Here, mini-batches are constructed by sampling pairs of step-diagram and part features, $({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{j})$ for optimizing the InfoNCE loss [11, 31]. Note that we randomly sample one part from their geometric-equivalent group. In this setup, pairs $({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{i})$ are considered as positive, otherwise negative. This loss encourages positive pairs to have higher similarity while negative pairs have lower similarity, allowing the similarity matrix to be used for generating the permutation matrix via Hungarian matching, as discussed in Sec. 3.3.
3.5.2 Losses for Pose Estimation.
Given a set of furniture parts, there may be subsets of parts that are geometrically identical to each other, such as the four legs of a table, the armrests on both sides of an armchair, or the support rods on the back of a chair. We denote geometrically equivalent parts as a part group $\mathcal{G}$. Within each group, we calculate the chamfer distance between these parts and the ground truth. Then, we apply Hungarian Matching (see Sec. 3.3) again with cost:
$\mathbf{C}_{ij}=\operatorname{CD}(\hat{R}_{i}\mathcal{P}_{i}+\hat{t}_{i},R_{j}\mathcal{P}_{j}+t_{j}),$ (9)
where CD denotes the chamfer distance measuring the similarity between unordered two point clouds [10], $i$ and $j$ index the parts within $\mathcal{G}$, and $\hat{R}$ and $\hat{t}$ denote the respective predicted rotation and translation. The optimal matching sequence is denoted as $\mathcal{M}$.
For the 3D part assembly task, following Li et al. [22], we supervise translation and rotation separately. We compute the $\ell_{2}$-distance for translation as follows:
$\mathcal{L}_{T}=\frac{1}{N}\sum_{i=1}^{N}\lVert\hat{t}_{[i]}-t_{i}\rVert_{2},$ (10)
where $[i]$ denotes the index of the matched part corresponding to the $i$-th part under optimal matching $\mathcal{M}$.
In addition to the geometric similarity between the parts mentioned above, each part itself may also have intrinsic symmetries. For instance, a cylindrical table leg has axial symmetry. Thus, we cannot simply use the absolute distance (e.g., $\ell_{2}$) on rotation as a supervision signal. Instead we use chamfer distance to supervise the rotation per part:
$\mathcal{L}_{C}=\frac{1}{N}\sum_{i=1}^{N}\operatorname{CD}(\hat{R}_{[i]}\mathcal{P}_{i},R_{i}\mathcal{P}_{i}).$ (11)
However, some parts may not be perfectly symmetric, e.g., due to small holes in different locations. To address these cases, we still encourage the model to penalize the $\ell_{2}$ distance for the rotation as a regularization term, formally expressed as:
$\mathcal{L}_{E}=\frac{1}{N}\sum_{i=1}^{N}\lVert\hat{R}_{[i]}\mathcal{P}_{i}-R_{i}\mathcal{P}_{i}\rVert_{2}.$ (12)
Additionally, to ensure that the predicted assembled shape closely matches the ground truth, we compute the chamfer distance between the predicted and ground truth assembled shapes:
$\mathcal{L}_{S}=\operatorname{CD}\left(\bigcup_{i=1}^{N}(\hat{R}_{[i]}\mathcal{P}_{i}+\hat{t}_{[i]}),\,\bigcup_{i=1}^{N}(R_{i}\mathcal{P}_{i}+t_{i})\right),$ (13)
where $\cup_{i=1}^{N}$ indicates the union of $N$ parts to form the assembled shape.
Finally, the overall loss for pose estimation is computed as a weighted sum of the aforementioned losses:
$\mathcal{L}_{\text{pose}}=\lambda_{T}\mathcal{L}_{T}+\lambda_{C}\mathcal{L}_{C}+\lambda_{E}\mathcal{L}_{E}+\lambda_{S}\mathcal{L}_{S}.$ (14)
4 Experiments
4.1 Datasets
Following [52, 55, 8, 56, 9, 27, 51, 22], we use PartNet [26], a comprehensive 3D shape dataset with fine-grained, hierarchical part segmentation, for both training and evaluation. In our work, we adopt the train/validation/test split used by Li et al. [22], where shapes with more than 20 parts are filtered out, and we use the finest granularity, Level-3, of PartNet. We focus on the two largest furniture categories: Chair and Table, containing 40,148 and 21,517 parts, respectively. Besides that, to demonstrate the generalizability of our method, we also evaluate its zero-shot capabilities on the IKEA-Manual [47] dataset, which features real-world IKEA furniture at the part level that are aligned with IKEA manuals. For consistency with PartNet, we evaluate both the Chair category, with 351 and 143 parts, and the Table, with 143 parts.
For each individual part, we sample 1,000 points from its CAD model and normalize them to a canonical space. Geometrically-equivalent part groups are identified based on the size of their Axis-Aligned Bounding Box (AABB). To generate the assembly manual, we render diagrammatic images of parts using Blender’s Freestyle functionality. More details of the preprocessing and rendering process are provided in the supplementary material.
[…]
Chair Table Chair Table Chair Table
[…]
Manual-PA w/o Order Manual 3.0 3.6 79.07 74.03 34.13 37.71
Manual-PA (Ours) Manual 1.7 1.8 89.06 87.41 58.03 56.66
[…]
Manual-PA w/o Order Manual 12.8 8.9 38.36 42.01 1.754 10.53
Manual-PA (Ours) Manual 11.4 4.8 42.51 49.72 3.509 15.79
Table 1: 3D part assembly results on the PartNet test split and Ikea-Manual. †: We re-trained the Image-PA model using diagrams (2D line drawing images) as the conditioning input instead of the original RGB images. The values in bold represent the best results, while underlined indicate the second.
4.2 Evaluation Metrics
To evaluate the effectiveness of 3D part assembly of different models, we employ three key metrics: Shape Chamfer Distance (SCD) [22], Part Accuracy (PA) [22], and Success Rate (SR) [23]. SCD quantifies the overall chamfer distance between predicted and ground truth shapes, providing a direct measure of assembly quality, scaled by a factor of $10^{3}$ for interpretability. PA assesses the correctness of individual part poses by determining whether the chamfer distance for a part falls below a threshold of $0.01$, reflecting the accuracy of each part’s alignment. Finally, SR evaluates whether all parts in an assembled shape are accurately predicted, offering a strict criterion for complete assembly success.
4.3 Main Results
4.3.1 Fully Supervised on PartNet.
In this setting, we train our model on PartNet for each category and test it on unseen shapes in the test split, which share similar patterns with the training set. As shown in Sec. 4.1, our method, Manual-PA, significantly outperforms all previous approaches in both the Chair and Table categories. Specifically, Manual-PA achieves superior performance compared to all other conditioned methods. Notably, when compared to Image-PA, we observe improvements of 26 and 17 part accuracy for the chair and table categories, respectively. This shows that for the 3D Part Assembly task, instruction manuals provide far better guidance than a single image or other conditioning inputs. “Manual-PA w/o Order” refers to our model without access to the part assembly order information, requiring it to learn this implicitly. Although the instruction manual provides strong visual guidance, performance improvements remain limited in this case, highlighting the challenges of implicit learning. By explicitly learning the correspondences between step diagrams and parts and applying positional encoding as soft guidance, we achieve PA improvements of approximately 10 and 13 in the chair and table categories, respectively. Note that all other methods, except for Image-PA and ours, are generative and follow the setting proposed by [52], where multiple assembly results are generated with varying Gaussian noise for the same set of parts. Only the most faithful assembly, as measured by Minimum Matching Distance (MMD) [1], is evaluated.
Figure 3: Comparison of average success rates (SR) across varying numbers of parts for different methods on PartNet chair test split. Number of chairs tested shown in background bar chart.
4.3.2 Number of Parts.
We analyze the assembly quality across varying numbers of parts on PartNet chair test split, as shown in Fig. 3. It is generally accepted that, assembling objects with more parts is more challenging due to the combinatorial explosive issue. However, our method consistently achieves a higher success rate across all part counts, demonstrating robust adaptability to different levels of assembly complexity. Notably, for shapes with more than 10 parts, 3DHPA and Image-PA exhibit near-zero success rates, whereas our method, Manual-PA, continues to produce competitive assembly results.
4.3.3 Zero-shot on IKEA-Manual.
We use the same model, pretrained on PartNet, to directly estimate the poses of furniture in IKEA-Manual, where all shapes are 3D models of real-world IKEA furniture, and each part corresponds to illustrations in their official assembly manuals. As shown in Sec. 4.1, our method, significantly outperforms strong competitors such as 3DHPA and Image-PA, demonstrating the strong generalizability of our approach to real world.
4.4 Ablation Studies
[…]
1 PE 4.3 72.22 22.50
2 PE ✓ 3.0 79.31 33.88
3 OneHot ✓ 3.0 79.07 34.13
4 OneHot ✓ ✓ 2.7 81.22 36.92
5 PE (Ours) ✓ ✓ 1.7 95.38 73.07
Table 2: Ablation results for various order designs on the PartNet chair test split. We use GT order for both training and inference.
[…]
Manual-PA (Ours) 0.789 1.7 89.06 58.03
 w/o batch-level sampling 0.410 2.3 56.93 15.68
 w/ Sinkhorn 0.774 1.7 88.25 56.26
 w/ GT Order 1.000 1.7 95.38 73.07
Table 3: Ablation results for various component designs in permutation learning on the PartNet chair test split. KT denotes the Kendall-tau coefficient that measure the ordinal correlation between predicted and GT order.
4.4.1 Effect of the Order for 3D Part Assembly.
We conduct experiments to explore various methods for incorporating order information in a 3D part assembly model, as shown in Sec. 4.4. First, by comparing lines 1 and 2, we observe that adding sequential information solely to part features improves assembly performance. We attribute this improvement to two factors: (1) distinguishing geometrically equivalent components, and (2) mitigating the combinatorial explosion problem, as claimed in SPAFormer [51]. Second, by comparing lines 2 and 5, where the latter provides both the step diagram and part features with correct correspondence, we achieve a significant performance boost. This indicates that explicitly establishing correspondence between steps and parts is beneficial, as learning these correspondences implicitly remains challenging for the model. Last, our empirical results show that, when position information is solely relevant to self-attention (lines 2 and 3), both PE and OneHot encodings yield similar performance. This observation aligns with prior transformer-based studies, which also utilize OneHot encodings [55, 9, 51]. However, in cases requiring cross-attention (line 4 and 5), PE outperforms OneHot encodings, highlighting its advantages in cross-attentive contexts.
4.4.2 Analysis of Permutation Learning Choices.
We conduct an ablation study for permutation learning, as shown in Sec. 4.4. First, we find that batch-level sampling is essential compared to applying contrastive learning individually within each manual book. This is primarily because contrastive learning benefits from a larger pool of negative samples. Second, while we employ the Sinkhorn [40] algorithm on the similarity matrix to enforce that rows and columns sum to one, empirical results indicate limited performance improvement. Last, by providing the model with the ground truth assembly order, we establish an upper performance bound for our assembly model.
Figure 4: Visualization of the cross attention scores between step diagrams (top) and a part (left, chair’s back) with PE at different positions. The final assembly results are displayed on the right.
Figure 5: Qualitative comparison of various 3D part assembly methods. Four examples are shown: chair (a) and table (b) from the PartNet dataset, and chair (c) and table (d) from the IKEA-Manual dataset.
4.5 Visualization and Analysis
4.5.1 Demonstration of Soft Guidance.
We manually add different positional encodings to the feature of chair’s back, as shown in Fig. 4. With PE(5), the model focuses on the final position, aligning well with the correct step diagram. However, when PE(2) is applied, drawing the model’s attention to the second step diagram, the model struggles to accurately predict the chair’s back pose. Notably, even with PE(0), the model manages to shift its attention toward the correct final step, demonstrating the fault tolerance and robustness of soft guidance: the model can still identify the correct match even when given an incorrect sequence.
4.5.2 Qualitative Comparison.
As shown in Fig. 5, our proposed Manual-PA produces the most faithful part pose prediction for the shapes illustrated in the step diagrams. For “Manual-PA w/o Order”, some parts do not align well with the assembly instructions, indicating the importance of providing precise correspondences between the step diagram and the parts for accurate assembly. In the case of Image-PA, the input consists only of the final image, making it challenging to locate the poses of parts that are occluded, such as the supporting rods beneath the table top in (b). 3DHPA lacks visual input and operates in a generative mode, relying solely on shape priors. Consequently, it generates shapes that are reasonable, as seen in (a), but there are noticeable discrepancies between its output and the ground truth shapes. For zero-shot results on the IKEA-Manual dataset, both 3DHPA and Image-PA exhibit degraded performance, while Manual-PA still produces faithful assembly results, demonstrating its superior generalization ability.
4.5.3 Cross Attention Map on Step Diagrams.
Figure 6: Visualization of cross-attention maps on step diagrams. The cross-attention represents the mean aggregation for each part across all step diagrams. Color red indicates a higher attention.
As shown in Fig. 6, we visualize the cross-attention maps between each part and the step diagrams. High attention values correspond to regions in the step diagram where the model pays more attention, which aligns with the spatial placement of parts during assembly. For instance, when the first blue board focuses on the back of the chair, it is positioned on the chair back, and when it attends to the area under the seat, it acts as a support between the legs. Comparing the two methods, we observe that without order as a soft guidance (represented by “Manual-PA w/o Order”), the attention regions are more dispersed, resulting in less accurate part placement. For example, the third cyan board does not focus entirely on the left chair leg without soft guidance, leading to a misaligned leg position. The absence of order as explicit guidance also results in incorrect part placement. For instance, the fourth purple board, which should be part of the seat, is instead assembled onto the chair back.
5 Conclusions and Future Work
We introduce a novel 3D part assembly framework that leverages diagrammatic manuals for guidance. Our proposed approach, Manual-PA, learns the sequential order of part assembly and uses it as a form of soft guidance for 3D part assembly. Manual-PA achieves superior performance on the PartNet dataset compared to existing methods and demonstrates strong generalizability to real-world IKEA furniture, as validated on the IKEA-Manual dataset.
Future work could aim to bridge the gap between synthetic and real-world IKEA manuals. First, the number of new parts introduced in each step can vary, requiring a more adaptable detection approach. Moreover, differing perspectives across steps necessitate robust viewpoint handling. The manuals also often include sub-module assemblies, resulting in a tree-like, non-linear assembly structure. Finally, while our current approach trains separate models for each furniture category, an ideal solution would be a unified model that generalizes across all kinds of furniture.