### v2

# Manual-PA: Learning 3D Part Assembly from Instruction Diagrams

Abstract
Assembling furniture amounts to solving the discrete-continuous optimization task of selecting the furniture parts to assemble and estimating their connecting poses in a physically realistic manner. The problem is hampered by its combinatorially large yet sparse solution space thus making learning to assemble a challenging task for current machine learning models. In this paper, we attempt to solve this task by leveraging the assembly instructions provided in diagrammatic manuals that typically accompany the furniture parts. Our key insight is to use the cues in these diagrams to split the problem into discrete and continuous phases. Specifically, we present Manual-PA, a transformer-based instruction Manual-guided 3D Part Assembly framework that learns to semantically align 3D parts with their illustrations in the manuals using a contrastive learning backbone towards predicting the assembly order and infers the 6D pose of each part via relating it to the final furniture depicted in the manual. To validate the efficacy of our method, we conduct experiments on the benchmark PartNet dataset. Our results show that using the diagrams and the order of the parts lead to significant improvements in assembly performance against the state of the art. Further, Manual-PA demonstrates strong generalization to real-world IKEA furniture assembly on the IKEA-Manual dataset.
1 Introduction
Figure 1: An illustration of the manual-guided 3D part assembly task. Given (a) a diagrammatic manual book demonstrating the step-by-step assembly process and (b) a set of texture-less furniture parts, the goal is to (c) infer the order of parts for the assembly from the manual sequence and predict the 6DoF pose for each part such that the spatially transformed parts assembles the furniture described in the manual.
With the rise of the DIY trend, nowadays it is the consumers who are increasingly assembling products like IKEA furniture [lee2021ikea, ben2021ikea, zhang2023aligning, zhang2024temporally, ben2024ikea], LEGO structures [wang2022translating, wan2024scanet], and exercise equipments [zhu2023multi, tian2022assemble]. This paradigm shift can reduce the product cost, but often poses challenges for untrained consumers, as an incorrect assembly in any step may lead to irreparable outcomes. Thus, automating the assembly or providing active help in the process is often desired.
Automated assembly faces the dual challenge of large combinatorial solution space, parameterized by the number of discrete parts in the assembly, and the need for a precise 6D pose estimation that correctly connects parts between each other. The limited number of feasible, stable assembly sequences makes this discrete-continuous optimization problem extremely challenging, even for advanced neural networks, especially as the number of parts increases [xu2024spaformer]. Fortunately, most assembly tasks include language-agnostic instruction manuals, such as diagrams, delineating the assembly. While such manuals help shrink the solution search space, the problem of correctly matching the 2D manual diagrams with their corresponding 3D object parts and inferring their assembly pose and order remains a significant challenge for automated systems.
This paper introduces a method to enable machines to learn how to assemble shapes by following a visual instruction manual, a task we call manual-guided 3D part assembly. As illustrated in Fig. 1, given a set of parts and a manual, our approach interprets the manual’s step-by-step diagrams to predict the precise 6D pose of each part, assembling them into the target shape. There are two lines of prior approaches that have similarities to our proposed task, namely methods that focus on the geometry of individual parts and inter-part relationships [li2020learning, zhan2020generative, zhang20223d, cheng2023score, zhang2024scalable, du2024generative], and methods such as MEPNet [wang2022translating] that learns the assembly in specific conditions, such as LEGO shapes. The former rely on priors like peg-hole joints [li2024category] or sequence generation [xu2024spaformer] to reduce search complexity but are generative and may yield unstable assemblies due to occlusions [li2020learning]. The latter often simplifies the task by assuming that parts are provided step-by-step, e.g., LEGO manuals. However, furniture manuals, such as IKEA, may not have such information, and it is necessary to determine which parts are used at each step before assembly. Additionally, while LEGO tasks feature standardized “stud” joints, general furniture assembly is more complex and lacks such constraints.
Based on the above observations, we identify two key challenges in our task. First, how to detect which part or parts are added at each step in the assembly diagrams. Given that these diagrams depict incremental assembly stages, the core challenge is to recognize newly added parts at each step. This requires aligning 2D assembly diagrams with the 3D parts to establish a correct assembly sequence. Second, how to effectively align the assembly process with the learned order. A straightforward approach is to directly use the learned sequence as the order for predicting poses, e.g., by adopting an auto-regressive prediction method similar to MEPNet [wang2022translating]. However, this approach carries the risk of accumulating errors—if a mistake occurs in an earlier step, subsequent steps are likely to fail as well. On the other hand, completely disregarding the sequence forces the assembly model to implicitly find correspondences, which makes both learning and prediction significantly more difficult. Therefore, the inferred sequence should serve as a soft guidance for the assembly process, helping each part focus more on the relevant step diagram.
To address these challenges, we present Manual-guided Part Assembly (Manual-PA), a novel transformer-based neural network that predicts correspondences between step diagrams and parts by computing their semantic similarity, which is learned using contrastive learning. Using this similarity matrix, we establish an assembly order via solving an optimal assignment problem between the aligned features of the 3D object parts and the ordered manual diagram steps, followed by permuting each part’s positional encoding according to the learned order. Finally, a transformer decoder fuses the two feature modalities to predict the final 6DoF pose for each part. Notably, the cross-attention between step diagrams and parts is guided by the positional encoding, ensuring that each step diagram receives a higher attention score for its corresponding part.
To empirically validate the effectiveness of Manual-PA, we conduct experiments on the standard PartNet benchmark dataset [mo2019partnet], on the categories of assembling the chair, table and storage classes. Our experiments and ablation studies clearly show that Manual-PA leads to state-of-the-art results on multiple metrics, including success rate, outperforming prior methods by a significant margin. Results further showcase the strong generalizability of our approach to real-world IKEA furniture assembly through experiments using the IKEA-Manual dataset [wang2022ikea].
We summarize our key contributions below:
  1. 1.
    We introduce a new problem setup of assembling 3D parts using diagrammatic manuals for guidance.
  2. 2.
    We present a novel framework, Manual-PA, for solving this task that learns the assembly order of the 3D parts, which is then used as soft guidance for the assembly.
  3. 3.
    We present experiments on the benchmark PartNet and IKEA-Manual datasets demonstrating state-of-the-art performances and cross-category generalization.
2 Related Works
2.0.1 3D Part Assembly
involves reconstructing complete shapes from sets of 3D parts, a task originally assisted by leveraging semantic part information [funkhouser2004modeling, sung2017complementme, yin2020coalesce, chaudhuri2011probabilistic, kalogerakis2012probabilistic, wu2020pq]. With the introduction of the PartNet dataset [mo2019partnet], recent works [zhan2020generative, narayan2022rgl, cheng2023score, zhang20223d, du2024generative, zhang2024scalable] leverage geometric features. For instance, 3DHPA [du2024generative] uses part symmetry for a hierarchical assembly, SPAFormer [xu2024spaformer] circumvents combinatorial explosion by generating a part order, and Joint-PA [li2024category] introduces a peg-hole constraint to facilitate assembly. However, these methods are generative towards producing plausible final shapes without aiming for preciseness. Approaches like GPAT [li2023rearrangement] and Image-PA [li2020learning] add guidance from point clouds or rendered images, respectively. Our work extends this direction by introducing guidance through instruction manuals, similar to how humans would do.
2.0.2 Permutation Learning
is often used in self-supervised methods for learning image representations [noroozi2016unsupervised, santa2017deeppermnet, kim2018learning, carlucci2019domain, pang2020solving]. However, in this work, we seek to align the 3D parts to instruction steps, for which we adopt a contrastive learning  [kim2021vilt, jia2021scaling, li2021align, radford2021learning] using a novel cross-modal alignment setup.
2.0.3 6DoF Object Pose Estimation
methods can be classified into instance-level [rad2017bb8, peng2019pvnet, xiang2017posecnn], category-level [tian2020shape, wang2019normalized], and unseen [labbe2022megapose, wen2024foundationpose]. Our approach relates to the unseen object category as furniture typically comprises novel parts. When the 3D models of these parts are available, many methods [chen2023zeropose, ornek2023foundpose, nguyen2024gigapose, lin2024sam, huang2024matchu] rely on feature matching followed by the PnP algorithm to infer pose. However, IKEA-style manuals are represented as line drawings making accurate feature extraction for matching to 3D parts challenging. Alternatively, some approaches [labbe2022megapose, wen2024foundationpose] use end-to-end models to directly predict the 6D pose. A common approach is to use CNOS [nguyen2023cnos] for object segmentation. However, in assembly diagrams, where multiple parts are connected and depicted together, SAM-based methods [kirillov2023segment] often struggle with accurate segmentation. Our approach addresses these challenges using an end-to-end framework that uses a cross-attention scheme to enable each part to attend to its relevant region in the step diagram.
Figure 2: Overview of our proposed method Manual-PA. (a) Feature extraction (Sec. 3.2): we extract semantic and geometrical features from both the step diagrams of the assembly manual and the corresponding part point clouds using the image encoder and point encoder, respectively. (b) Manual-guided part permutation learning (Sec. 3.3): we compute a similarity matrix $\mathbf{S}$ between the two modalities, and subsequently apply the Hungarian algorithm to obtain the permutation matrix $\mathbf{P}$ for the parts. (c) Manual-guided part pose estimation (Sec. 3.4): we add positional encodings (PE) $\Phi$ to both the step diagrams and parts, where the PE order for parts is determined by the order predictions from (b). This is followed by a transformer decoder to enable multimodal feature fusion and interaction, along with a pose prediction head to determine the rotation $R_{i}$ and translation $t_{i}$ of each part. The predicted poses are then applied to the corresponding parts to obtain the final assembled shape $\mathbf{S}$.
3 Manual-Guided Part Assembly
In this section, we first provide a concrete formal definition of the manual-based 3D part assembly task in Sec. 3.1. We then present a detailed explanation in Secs. 3.2, 3.3 and 3.4 of the proposed learning pipeline as depicted in Fig. 2, followed by a description of the training process and the loss functions used for each stage in Sec. 3.5.
3.1 Problem Formulation
Our task involves two different input modalities: a set of 3D parts and a corresponding instruction manual. The 3D parts are represented as a multiset [^1] We allow multiple identical parts, e.g., chair legs, to appear in the set. of unordered point clouds, denoted by $\{\mathcal{P}_{i}\}_{i=1}^{N}$. Note that the number of parts $N$ may differ from furniture to furniture. The manual consists of an ordered sequence of $N$ step diagram images, denoted by $(\mathcal{I}_{1},\mathcal{I}_{2},\ldots,\mathcal{I}_{N})$; each diagram depicting an incremental step in the assembly process of a 3D furniture shape $\mathcal{S}$. These step diagrams are 2D line drawings, which are 2D projections of textureless 3D CAD models of these parts captured from a camera $\mathcal{C}$, and we assume for each step, only one additional part is attached to the furniture assembled thus far. Our goal is to align the 3D parts with their delineations in the respective manual steps towards predicting the assembly order of the parts and the part poses $\{(R_{i},t_{i})\in SE(3)\}_{i=1}^{N}$ for the final assembled furniture, where $R_{i},t_{i}$ represent the $3\times 3$ rotation matrix and the three-dimensional translation vector respectively for the $i$-th part in the SE(3) transformation space. That is, these parts undergo rigid transformations such that their combination constructs the complete 3D furniture shape $\mathcal{S}$ as seen from the viewpoint of camera $\mathcal{C}$.
3.2 Feature Extraction
We first use off-the-shelf encoders to extract semantic latent features from the inputs. For the 3D part point clouds $\{\mathcal{P}_{i}\}_{i=1}^{N}$, we utilize a lightweight version of PointNet [qi2017pointnet], similar to the one used in Image-PA [li2020learning], followed by a linear layer to map the features to a common dimension $D$, resulting in geometric part features ${\mathbf{f}^{\text{P}}}\in\mathbb{R}^{N\times D}$. For sequences of step diagrams $(\mathcal{I}_{j})_{j=1}^{N}$, given that the assembly instructions are provided step-by-step in an incremental manner, our primary focus is on the differences between consecutive steps. For any two consecutive diagram images $\mathcal{I}_{j}$ and $\mathcal{I}_{j+1}$, for $j=1,2,\ldots,N-1$, we compute the difference image given by $|\mathcal{I}_{j}-\mathcal{I}_{j+1}|$ that shows which new part is added to the thus-far assembled furniture. Next, this difference image is patchified into $K$ image patches, that are then fed to DINOv2 image encoder [caron2021emerging, oquab2023dinov2] to extract features, followed by a linear layer to map these features to the same dimensionality $D$, resulting in semantic image features ${\mathbf{f}^{\text{I}}}\in\mathbb{R}^{N\times K\times D}$.
3.3 Manual-Guided Part Permutation Learning
In our task, we seek to learn the order of a set of parts conditioned on an ordered sequence of step diagrams. This problem is similar to learning the feature correspondences between two modalities in order to align them, with an ideal alignment corresponding to a permutation matrix. Specifically, our objective is to derive a similarity matrix between the two modalities such that each modality feature is unambiguously assigned to its corresponding feature in the other modality and we desire this assignment matches with the ground truth order in the step sequence in the manual. Formally, we construct the similarity matrix $\mathbf{S}\in\mathbb{R}^{N\times N}$ as:
$\mathbf{S}_{ij}=\operatorname{sim}({\mathbf{f}_{i}^{\text{P}}},\textbf{g}^{\text{I}}_{j}),$ (1)
where $\textbf{g}^{\text{I}}\in\mathbb{R}^{N\times D}$ is derived from the $N\times K\times D$-dimensional step-diagram features $\mathbf{f}^{\text{I}}$} by performing a max-pooling operation on the patch dimension $K$. We measure the similarity in the feature space using the dot product as $\operatorname{sim}(\mathbf{a},\mathbf{b})=\mathbf{a}\cdot\mathbf{b}$.
Next, we obtain the permutation matrix $\mathbf{P}\in\{0,1\}^{N\times N}$ by solving a bipartite matching optimization problem using the Hungarian matching algorithm [kuhn1955hungarian] with cost $\mathbf{C}=-\mathbf{S}$:
$\begin{array}[]{rl}\text{minimize}&\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbf{C}_{ij}\mathbf{P}_{ij}\\ \text{subject to}&\sum_{j=1}^{N}P_{ij}=1,\quad\forall i\in\{1,\dots,N\}\\ &\sum_{i=1}^{N}P_{ij}=1,\quad\forall j\in\{1,\dots,N\}\\ &P_{ij}\in\{0,1\},\quad\forall i,j\in\{1,\dots,N\},\end{array}$ (6)
where we minimize the cost, subject to constraints ensuring $\mathbf{P}$ is a permutation matrix, i.e., each row and column of $\mathbf{P}$ contains exactly one entry equal to $1$, with all other entries being 0. The final part order can be obtained by $\sigma=\operatorname{argmax}(\mathbf{P})$ where $\operatorname{argmax}$ is applied columnwise.
3.4 Manual-Guided Part Pose Estimation
3.4.1 Positional Encoding.
Since the attention mechanism in the transformer is permutation-invariant, we follow vaswani2017attention [vaswani2017attention] in using positional encoding to convey the order of both step diagrams and parts to the model. Formally, the positional encoding is defined as a function that maps a scalar to a sinusoidal encoding, $\phi:\mathbb{N}\rightarrow\mathbb{R}^{D}$:
$\phi(x)_{2i}=\sin\left(\frac{x}{\tau_{p}^{2i/d}}\right)\!,\;\phi(x)_{2i-1}=\cos\left(\frac{x}{\tau_{p}^{2i/d}}\right)\!,$ (7)
where $i=1,\dots,d/2$ indexes the dimension and $\tau_{p}$ is a temperature hyperparameter. By applying $\phi$ to encode the sequential order of step diagrams as they appear in the manual (i.e., using $x$ as a simple increasing sequence $1,2,\dots,N$), we obtain a positional encoding $\Phi\in\mathbb{R}^{N\times D}$, where ${\mathbf{p}^{\text{I}}}:=\Phi$. Given the permutation matrix $\mathbf{P}$ that corresponds to the order of the parts, we can compute the positional encoding for the parts as ${\mathbf{p}^{\text{P}}}=\mathbf{P}^{T}\Phi$. During training, we use the ground truth part order, whereas at inference time, we replace it with the predicted permutation matrix $\hat{\mathbf{P}}$, yielding $\hat{\mathbf{p}}^{\text{P}}=\hat{\mathbf{P}}^{T}\Phi$. These positional encodings are then added to the respective features before being fed into the transformer decoder and provide soft guidance on the sequence order. We adopt RoPE [su2024roformer] for better performance.
3.4.2 Transformer Decoder.
The modality interaction and fusion between step diagrams and parts occur through a stack of $L$ transformer [vaswani2017attention] decoder layers. In each layer, the self-attention module receives the output from the previous layer (initially, the part features ${\mathbf{f}^{\text{P}}}+{\mathbf{p}^{\text{P}}}$ in the first layer) to perform part-to-part interactions. This is followed by a cross-attention module that injects and fuses the step diagram features ${\mathbf{f}^{\text{I}}}+{\mathbf{p}^{\text{I}}}$ with the part features processed by the self-attention module, enabling step-to-part interactions. Notably, for the step diagram features, we concatenate the features from all diagrams from the same manual along the patch dimension, resulting in a tensor of shape $\mathbb{R}^{NK\times D}$.
3.4.3 Pose Prediction Head.
Finally, the output $O_{\text{Dec}}$ from the transformer decoder is fed into a pose prediction head, which predicts the translation and rotation for each part, denoted as $(\hat{R}_{i},\hat{t}_{i})$. In our implementation, the rotation matrix $\hat{R}_{i}$ is represented as a unit quaternion $\hat{q}_{i}\in\mathbb{R}^{4}$. We initialize its bias to an identity quaternion $(1,0,0,0)$ and normalize the prediction to ensure a unit norm.
3.5 Training and Losses
First, we train the manual-guided part order learning model until it successfully aligns the two modalities and provides an optimal permutation for the parts. Subsequently, we train the transformer decoder-based part pose estimation model using the predicted order from the first model to infer the part poses. Our empirical results indicate that having an accurate part order is crucial for enhancing the inference performance of the pose estimation network.
3.5.1 Loss for Permutation Learning.
Borrowing the notation from Secs. 3.2 and 3.3, we define the contrastive-based order loss as:
$\mathcal{L}_{\text{order}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\operatorname{sim}({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{i})/\tau)}{\sum_{j=1}^{B}\exp(\operatorname{sim}({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{j})/\tau)},$ (8)
where $B$ is the mini-batch size, $\tau$ is the temperature, and $\sigma(i)$ denotes the index for the $\sigma(i)$-th part that corresponds to the $j$-th step according to the ground truth part order $\sigma$. Here, mini-batches are constructed by sampling pairs of step-diagram and part features, $({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{j})$ for optimizing the InfoNCE loss [hadsell2006dimensionality, oord2018representation]. Note that we randomly sample one part from their geometric-equivalent group. In this setup, pairs $({\mathbf{f}^{\text{P}}}_{\!\!\sigma(i)},\textbf{g}^{\text{I}}_{i})$ are considered as positive, otherwise negative. This loss encourages positive pairs to have higher similarity while negative pairs have lower similarity, allowing the similarity matrix to be used for generating the permutation matrix via Hungarian matching, as discussed in Sec. 3.3.
3.5.2 Losses for Pose Estimation.
Given a set of furniture parts, there may be subsets of parts that are geometrically identical to each other, such as the four legs of a table, the armrests on both sides of an armchair, or the support rods on the back of a chair. We denote geometrically equivalent parts as a part group $\mathcal{G}$. Within each group, we calculate the chamfer distance between these parts and the ground truth. Then, we apply Hungarian Matching (see Sec. 3.3) again with cost:
$\mathbf{C}_{ij}=\operatorname{CD}(\hat{R}_{i}\mathcal{P}_{i}+\hat{t}_{i},R_{j}\mathcal{P}_{j}+t_{j}),$ (9)
where CD denotes the chamfer distance measuring the similarity between unordered two point clouds [fan2017point], $i$ and $j$ index the parts within $\mathcal{G}$, and $\hat{R}$ and $\hat{t}$ denote the respective predicted rotation and translation. The optimal matching sequence is denoted as $\mathcal{M}$.
For the 3D part assembly task, following li2020learning [li2020learning], we supervise translation and rotation separately. We compute the $\ell_{2}$-distance for translation as follows:
$\mathcal{L}_{T}=\frac{1}{N}\sum_{i=1}^{N}\lVert\hat{t}_{[i]}-t_{i}\rVert_{2},$ (10)
where $[i]$ denotes the index of the matched part corresponding to the $i$-th part under optimal matching $\mathcal{M}$.
In addition to the geometric similarity between the parts mentioned above, each part itself may also have intrinsic symmetries. For instance, a cylindrical table leg has axial symmetry. Thus, we cannot simply use the absolute distance (e.g., $\ell_{2}$) on rotation as a supervision signal. Instead we use chamfer distance to supervise the rotation per part:
$\mathcal{L}_{C}=\frac{1}{N}\sum_{i=1}^{N}\operatorname{CD}(\hat{R}_{[i]}\mathcal{P}_{i},R_{i}\mathcal{P}_{i}).$ (11)
However, some parts may not be perfectly symmetric, e.g., due to small holes in different locations. To address these cases, we still encourage the model to penalize the $\ell_{2}$ distance for the rotation as a regularization term, formally expressed as:
$\mathcal{L}_{E}=\frac{1}{N}\sum_{i=1}^{N}\lVert\hat{R}_{[i]}\mathcal{P}_{i}-R_{i}\mathcal{P}_{i}\rVert_{2}.$ (12)
Additionally, to ensure that the predicted assembled shape closely matches the ground truth, we compute the chamfer distance between the predicted and ground truth assembled shapes:
$\mathcal{L}_{S}=\operatorname{CD}\left(\bigcup_{i=1}^{N}(\hat{R}_{[i]}\mathcal{P}_{i}+\hat{t}_{[i]}),\,\bigcup_{i=1}^{N}(R_{i}\mathcal{P}_{i}+t_{i})\right),$ (13)
where $\cup_{i=1}^{N}$ indicates the union of $N$ parts to form the assembled shape.
Finally, the overall loss for pose estimation is computed as a weighted sum of the aforementioned losses:
$\mathcal{L}_{\text{pose}}=\lambda_{T}\mathcal{L}_{T}+\lambda_{C}\mathcal{L}_{C}+\lambda_{E}\mathcal{L}_{E}+\lambda_{S}\mathcal{L}_{S}.$ (14)
4 Experiments
4.1 Datasets
Following [zhan2020generative, zhang20223d, cheng2023score, zhang2024scalable, du2024generative, narayan2022rgl, xu2024spaformer, li2020learning], we use PartNet [mo2019partnet], a comprehensive 3D shape dataset with fine-grained, hierarchical part segmentation, for both training and evaluation. In our work, we adopt the train/validation/test split used by li2020learning [li2020learning], where shapes with more than 20 parts are filtered out, and we use the finest granularity, Level-3, of PartNet. We focus on the three largest furniture categories: chair, table and storage, containing 40,148, 21,517 and 9,258 parts, respectively. Besides that, to demonstrate the generalizability of our method, we also evaluate its zero-shot capabilities on the IKEA-Manual [wang2022ikea] dataset, which features real-world IKEA furniture at the part level that are aligned with IKEA manuals. For consistency with PartNet, we evaluate both chair category with 351 parts, and table, with 143 parts. Storage is excluded due to only 3 samples.
For each individual part, we sample 1,000 points from its CAD model and normalize them to a canonical space. Geometrically-equivalent part groups are identified based on the size of their Axis-Aligned Bounding Box (AABB). To generate the assembly manual, we render diagrammatic images of parts using Blender’s Freestyle functionality. More details of the preprocessing and rendering process are provided in the supplementary material.
Table 1: 3D part assembly results on the PartNet test split and Ikea-Manual. †: We re-trained the Image-PA model using diagrams (2D line drawing images) as the conditioning input instead of the original RGB images. The values in bold represent the best results, while underlined indicate the second. ‡: Average over five runs.
\__nicematrix_patch_booktabs:\__nicematrix_revert_colortbl: